Breeding rabbits--please determine the relative po
Answer #1

1). In the trihybrid cross, the offspring with the least number represent the double recombinant ones, and the offspring with the highest number represent the parental genotypes, the remaining is monohybrid cross.

Assume that the gene coding for black is B (b codes for white), gene coding for two-headed is H (h codes for one-head), gene coding for long eared is L (l codes for short eared).

260 black, two-headed, long eared >Parental 280 white, one-headed, short eared 16 white, two-headed, long eared 14 black, one

From the double crossover and single crossover gametes, we can say that the gene order is HBL.

Now, distance between H and B is,

= number of gametes produced by single crossover of H and B + number of double crossover gametes*100/ total number of gametes.

= (130+150+16+14)*100/ 1000 = 31 map units (means, 31% recombination).

Now, distance between B and L

= number of gametes produced by single crossover of B and L + number of double crossover gametes*100/ total number of gametes.

= (70+80+16+14)*100/ 1000 = 18 map units (means, 18% recombination).